Circular No. 6892

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J0840.7+2248 J. Heise and J. in ’t Zand, Space Research Organization of the Nether-

lands, Utrecht; R. Ricci, G. d’Andreta, and J. Muller, BeppoSAX, Rome; and L. Piro, Istituto di Astrophysica Spaziale, Rome, report: “A bright, high-latitude fast x-ray transient source has been observed on Apr. 29.366 UT in one of the Wide Field Cameras of BeppoSAX. The event lasted ∼ 110 s. The peak intensity of the structured event reached ∼ 1 Crab (2–25 keV). Its position is determined at α = 8^{h}40^{m}40^{s}, δ = +22^{o}48′.3 (equinox 2000.0; error radius 3′) and the source is designated SAX J0840.7+2248. A follow-up observation with the BeppoSAX narrow-field instruments is in progress.”

MARKARIAN 421 P. Grandi, Istituto di Astrofisica Spaziale, CNR, Rome; and L. Chi-

appetti, Istituto di Fisica Cosmica e Tecnologie Relative, CNR, Milan; L. Maraschi, Osservatorio Astronomico di Brera, Milan; and E. Pian, Istituto di Tecnologie e Studio delle Radiazioni Extraterrestri, CNR, Bologna; on behalf of the BeppoSAX collaboration on the coordinated campaign for spectral variability of Mkn 421, report: “This object was observed with the BeppoSAX Narrow Field Instruments on Apr. 21.08–22.13 UT. The source was bright and variable, and it showed a well-defined flare with an initial 50-percent increase over 15 000 s, followed by a slower regular decrease to about half the peak intensity. The source was well visible in the Phoswich Detector System up to ∼ 50 keV. The preliminary standard analysis indicates an average energy flux of 3.2 × 10^{−10} erg cm^{−2} s^{−1} in the band 2–10 keV, close to the maximum reported by ASCA on Apr. 27 (IAUC 6888). Another BeppoSAX observation on Apr. 23–24 found an average flux of 1.9 × 10^{−10} erg cm^{−2} s^{−1} (2–10 keV). Thus the bright active state extends at least for a week and may persist for a longer time. Continuing observations at all wavelengths are urged.”

SUPERNOVA 1998bp IN NGC 6495 S. Nakano, Sumoto, Japan, forwards the following position and mag-

nitude for SN 1998bp (cf. IAUC 6890) from a CCD image obtained with a 0.4-m f/6 reflector by R. Kushida at Yatsugatake South Base Observatory on Apr. 30.75 UT: α = 17^{h}54^{m}50^{s}.74, δ = +18^{o}19′50′′.5 (equinox 2000.0), V = 15.0 (measurer Y. Kushida).

1998 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green